Sodium cyanide is the salt of the weak acid HCN. Calculate the concentration of H3O+ ; OH- ; HCN and Na+ in solution prepared by dissolving 10.8 g of NaCN in enough water to make 5.00 x 102 mL of solution at 25 0C.
1
Expert's answer
2020-11-17T10:36:47-0500

"M(NaCN)= 49 g\/mol" (ref.1)

The overall volume "V = 0.51 L"

"n(NaCN) = m(NaCN)\/M(NaCN) ="

"= 10.8 g/49(g/mol) \approx 0.22041 mol"

"C(NaCN) = n(NaCN)\/V ="

"= 0.22041mol / 0.51L \approx 0.4322mol/L"

"NaCN" is a strong electrolyte, therefore assume it dissociates completely:

"NaCN \to Na^+ (aq.) + CN^-(aq.)"

Therefore, "C(NaCN) = C(Na^+(aq.)) ="

"= C(CN^-(aq.)) = 0.4322mol\/L"

Cyanide ion undergoes hydrolysis:

"CN^-(aq.) + H_2O \leftrightarrows HCN + OH^-(aq.)"

The dissociation of water is ignored

"C_{reacted}(CN^-(aq.)) = [NCN] = [OH^-(aq.)]="

"= x\ mol/L"

"[CN^-(aq.)]="

"= C(CN^-(aq.)) - C_{reacted}(CN^-(aq.)) ="

"= 0.4322mol\/L - x"

"K_h = [HCN][OH^-(aq.)]\/[CN^-] ="

"= x^2\/(0.4322 - x) = 2.5*10^{-5}" (ref. 2)

"x^2 = 2.5*10^{-5} *(0.4322 - x)"

"x^2 = 1.0805*10^{-5} - 2.5*10^{-5}x"

"x^2 + 2.5*10^{-5}x - 1.0805*10^{-5} = 0"

Solving the equation gives "x \approx 0.00327462 \ mol/L = 3.27462 *10^{-3} \ mol/L"

"[HCN] = [OH^-(aq.)] = 3.27462 *10^{-3} \ mol/L"

at "t = 25 \ ^\omicron C" the following holds:

"pH + pOH = 14"

"pH = 14 - pOH"

"pOH = - log[OH^-(aq.)] ="

"= - log(3.27462 *10^{-3}) \approx 2.48484"

"pH = 14 - 2.48484 = 11.51516"

"pH = - log[H_3O^+(aq.)]"

"[H_3O^+(aq.)] = 10^{-pH} ="

"=10^{-11.51516} mol/L \approx 3.0538*10^{-12} \ mol/L"

"[Na^+(aq.)] = C(Na^+(aq.)) = 0.4322\ mol/L"



Answer:

"[H_3O^+(aq.)] = 3.0538*10^{-12} \ mol/L;"

"[HCN] = [OH^-(aq.)] = 3.27462 *10^{-3} \ mol/L;"

"[Na^+(aq.)] = 0.4322\ mol/L."


References:

1. https://pubchem.ncbi.nlm.nih.gov/compound/Sodium-cyanide#section=Computed-Properties

2. https://pubchem.ncbi.nlm.nih.gov/compound/Sodium-cyanide#section=Stability-Shelf-Life